OpenStudy (anonymous):

A 26 foot long ladder is leaning against a vertical wall. The foot of the ladder is 10 feet away from the base of the wall. The foot of the ladder is being pulled away from the base of the wall at a rate of 4 feet per second. How fast is the top of the ladder sliding down the wall at this instant? (related rates)

myininaya (myininaya):

related rates problem awesome!

myininaya (myininaya):

cyter beat me

myininaya (myininaya):

lol wrong problem

OpenStudy (anonymous):

could both answers be correct?

myininaya (myininaya):

lol there are two different answers let me see if i made a mistake

OpenStudy (anonymous):

You used 4 instead of 10

myininaya (myininaya):

cyter is right

myininaya (myininaya):

i used y' for y my bad

myininaya (myininaya):

cyter wins

OpenStudy (anonymous):

we both win because we spent energy working on the problem

myininaya (myininaya):

lol

OpenStudy (anonymous):

in the writing in the purple, isn't the derivative of s^2 2s, not just 2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

the right hand side should have been 2s ds/dt I knew something was amiss.

myininaya (myininaya):

since s is a constant (it doesnt change) s'=0 so 2ss'=0

myininaya (myininaya):

so it doesnt change the answer

OpenStudy (anonymous):

yep it makes no difference but it looks logical

myininaya (myininaya):

right and it is best to be logical during to look crazy

myininaya (myininaya):

than not during*

OpenStudy (anonymous):

but what happens to the 2s? does it cancel out? and how does the x (dx/dt) becomes negative?

myininaya (myininaya):

2ss'=0

myininaya (myininaya):

the derivative of s^2 is 2ss' but since s doesnt change s'=0

OpenStudy (anonymous):

oo right, sorry

OpenStudy (anonymous):

factor out the 2 s and devide both sides by 2 0/2 = 0

myininaya (myininaya):

cyter did you draw that on the computer? it looks very mechanical like

myininaya (myininaya):

on some parts

myininaya (myininaya):

lol looks better than my horrible hand writing

OpenStudy (anonymous):

\[\text{2((xdx/dt)+(ydy/dt) = 2sds = 0/dt}\] dividing both sides by 2 we have \[\text{((xdx/dt)+(ydy/dt) = sds/dt = 0}\] now subtract both sides by (xdx/dt) we get \[\text{(ydy/dt) = -(xdx/dt)}\] and then divide both sides by y to get dy/dt alone \[\text{(dy/dt) = -(x/y)dx/dt}\] Then plug and play

OpenStudy (anonymous):

yes I used paint which isn't very good

OpenStudy (anonymous):

Your handwriting is legible, unlike mine where I have to use a drawing program on my computer

myininaya (myininaya):

i have a scanner haha

myininaya (myininaya):

lol

OpenStudy (anonymous):

ok i get it now, thank you both of you for your help

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

Using cyter's solution diagram, \[y=x \text{Tan}[\text{ArcCos}[x/26]] \] then \[y=26 \sqrt{1-\frac{x^2}{676}} \] The total derivative of the above is: \[\text{Dt}[y]==-\frac{x \text{Dt}[x]}{26 \sqrt{1-\frac{x^2}{676}}} \] Replace x with 10 and Dt[x] with 4 and simplify, \[\text{Dt}[y]==-\frac{5}{3} \] I hope there are no errors. First DE solved in years.